OpenStudy (jonnywktn):

can someone help me with the following Fourier series question F(x) = {-9 when - π

OpenStudy (jonnywktn):

i have worked out the fourier series i just need to work out b for n = 1,2,3,4,5 if this makes any more sense

OpenStudy (jonnywktn):

happy to pay for the correct answers

OpenStudy (holsteremission):

I'm afraid that doesn't make any more sense. Are you trying to work out the coefficients of the sine series?

OpenStudy (holsteremission):

Your function is odd, so I'm assuming that's the case, as both \(a_0\) and \(a_n\) are zero for all \(n\in\mathbb{N}\). You have \[b_n=\frac{1}{\pi}\int_{-\pi}^\pi f(x)\sin(nx)\,\mathrm{d}x=\frac{9}{\pi}\left(\int_0^\pi\sin(nx)\,\mathrm{d}x-\int_{-\pi}^0\sin(nx)\,\mathrm{d}x\right)\]Are you having trouble computing these integrals?

OpenStudy (holsteremission):

Replacing \(x\mapsto -x\), you can consolidate the integral to get \[b_n=\frac{18}{\pi}\int_0^\pi \sin(nx)\,\mathrm{d}x\]Surely you know the chain rule?

OpenStudy (jonnywktn):

i have bn = 18/nπ [1-(-1)n] just been asked to work out b for 1,2,3,4,5 and really cant remember how to do it

OpenStudy (jonnywktn):

b for n = 1,2,3,4,5 if you could show me how to do them 5 it would be much appreciated

OpenStudy (holsteremission):

That's just a matter of replacing \(n\) with each of those values. For instance, \[b_1=\frac{18}{1\times\pi}(1-(-1)^1)=\frac{18}{\pi}(1+1)=\frac{36}{\pi}\]

OpenStudy (jonnywktn):

so what would b 2 and 3 be? just so im clear

OpenStudy (holsteremission):

For even values of \(n\), notice that \((-1)^n=1\), so both \(b_2\) and \(b_4\) would be \(0\), since \[\frac{18}{n\pi}(1-1)=0\] When \(n\) is odd, \((-1)^n=-1\). So for instance, when \(n=3\), you get \[\frac{18}{3\pi}(1-(-1)^3)=\frac{6}{\pi}(1+1)=\frac{12}{\pi}\]

OpenStudy (jonnywktn):

so for b2 its 18/2π (1-1) = 0

OpenStudy (jonnywktn):

and b5 would be 18/5π (1-(-1)5) = 3.6/π (1+1) = 7.2/π ??? is this correct

OpenStudy (jonnywktn):

????

OpenStudy (holsteremission):

That's correct, yes